What is the annual worth for option A if the initial cost is $100 and there is a uniform annual benefit of $10 for an infinite time frame. Interest rate is 8%. Do not put $ in answer.

Hint: Use equation for infinite time frame: A=P*I and remember cost is negative and that you also have an annual benefit to add to the annual cost associated with this option.
Answer #1

Benefit from the option A=10 lifetime=10+10/1.08 + 10/1.082 +.....

Benefit=10/1-(1/1.08)=10*1.08/0.08=135

And cost= 100

Thus new worth of the option A=135-100=35